Exercise 15-20 Presented below is information from the annual report of Cullumber Plastics, Inc. Operating income $515,600 Bond interest expense 123,000 392,600 Income taxes 165,176 Net income $227,424 Bonds payable $1,000,000 Common stock 897,000 Retained earnings 339,000 (a) Compute the return on common stockholders’ equity and the rate of interest paid on bonds. (Assume balances for debt and equity accounts approximate averages for the year.) (Round answers to 1 decimal place, e.g. 16.8%.) Rate of return on common stock equity % Rate of interest paid on bonds payable % (b) Is Cullumber Plastics, Inc. trading on the equity successfully?
Answer #1

Answer:

Return on Common Stockholders Equity = Net Income / ( Common stock + Retained earnings )

                                                           = 227424 / (897000 + 339000 )

                                                           = 227424 / 618000

                                                           = 36.8%

Rate of Interest on Bonds Payable = Bond interest expense / Bonds payable = 123000 / 1000000 = 12.3%

The return on common Stockholder’s equity is greater than the rate of interest paid on bonds. The firm is trading on the equity successfully.